# The complex structure of the Teichmüller space of circle diffeomorphisms in the Zygmund smooth class

### v2

@@ 4. The little Zygmund smooth class
In this section, we supply additional results on the little class of the Zygmund smoothness. For the space ${\rm Bel}^{Z}(\mathbb{D}^{*})$ of the Beltrami coefficients, its little subspace is defined by
${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})=\{\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})\mid\lim_{t\to 0}\ {\rm ess}\!\!\!\!\!\!\!\!\!\!\sup_{1<|z|<1+t\quad\ }\!\!\!\!\!\!(|z|^{2}-1)^{-1}|\mu(z)|=0\}.$
The Teichmüller space of the little Zygmund smooth class is defined by $T^{Z}_{0}=\pi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$. We remark that the usage of naught ${}_{0}$ in this section is different from [9, 14].
We see that ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ is a closed subspace of ${\rm Bel}^{Z}(\mathbb{D}^{*})$. Indeed, assuming that a sequence $\{\mu_{k}\}_{k\in\mathbb{N}}$ in ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ and $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ are given so that $\|\mu_{k}-\mu\|_{Z}\to 0$ as $k\to\infty$, we show that $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. For each $\varepsilon>0$, we can choose some $k_{0}\in\mathbb{N}$ such that $\|\mu_{k_{0}}-\mu\|_{Z}<\varepsilon$. Since $\mu_{k_{0}}\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, there exists some $t_{0}>0$ such that
${\rm ess}\!\!\!\!\!\!\!\!\!\!\!\sup_{1<|z|<1+t_{0}\quad\ }\!\!\!\!\!\!\!(|z|^{2}-1)^{-1}|\mu_{k_{0}}(z)|<\varepsilon.$
@@
uniformly as $t\to 0$. We can show that $(f^{\mu})^{\prime}$ satisfies the little Zygmund condition if and only if $\mu$ belongs to ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. Indeed, this is verifed by examining the proof of [13, Theorem 1.1] with the fact that $(f_{\mu})^{\prime}|_{\mathbb{S}}$ satisfies the little Zygmund condition similar to the case of self-diffeomorphisms of $\mathbb{S}$ if and only if
@@
To see that $\Phi:{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})\to A^{Z}_{0}(\mathbb{D})$ is a holomorphic split submersion, we verify that for every $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, the local holomorphic map $s_{\mu}:V_{\varphi}\to{\rm Bel}^{Z}(\mathbb{D}^{*})$ given in Theorem 7 maps $A^{Z}_{0}(\mathbb{D})\cap V_{\varphi}$ into ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. We first choose $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric on $\mathbb{D}^{*}$. By examining the proof of Lemma 6, we can choose such $\nu$ in ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. Then, as in the proof of [9, Lemma 7.5], we have
@@
for every $\varphi^{\prime}\in V_{\varphi}$ and every $z^{*}=1/\bar{z}\in\mathbb{D}^{*}$. If $\varphi^{\prime}\in A^{Z}_{0}(\mathbb{D})$ in addition, then $\varphi^{\prime}-\varphi\in A^{Z}_{0}(\mathbb{D})$ and $\nu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ imply $s_{\nu}(\varphi^{\prime})\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$.
For the arbitrary $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, we apply the right translations $r_{\mu}$ and $r_{\nu}$, which are biholomorphic automorphisms of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ by Lemma 5, and in fact of ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ as we can see by the proof of Proposition 4. Then, $s_{\mu}=r_{\mu}^{-1}\circ r_{\nu}\circ s_{\nu}$ is a local holomorphic right inverse of $\Phi$ defined on $V_{\varphi}\cap A^{Z}_{0}(\mathbb{D})$ satisfying $s_{\mu}\circ\Phi(\mu)=\mu$. ∎
Corollary 15.
@@
The claim $\Phi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))=A^{Z}_{0}(\mathbb{D})\cap\Phi({\rm Bel}(\mathbb{D}^{*}))$ is seen by the proof of Theorem 14 because $s_{\mu}(A^{Z}_{0}(\mathbb{D})\cap V_{\varphi})\subset{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ implies $A^{Z}_{0}(\mathbb{D})\cap V_{\varphi}\subset\Phi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$. The other claims directly follows from Theorem 14. ∎
@@
The former claim is proved similarly to the case of the Schwarzian derivative map $\Phi$ as in the proof of Theorem 14. For the latter claim, we use the holomorphic map $\Lambda:B^{Z}(\mathbb{D})\to A^{Z}(\mathbb{D})$ with $\Lambda\circ\Psi=\Phi$, which is handled in Lemma 12. Since $\Phi$ is a holomorphic split submersion on ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ by Theorem 14, we see that so is $\Psi$. ∎

### v3

@@ 4. The little Zygmund smooth class
In this section, we supply additional results concerning the little class of Zygmund smoothness. For the space ${\rm Bel}^{Z}(\mathbb{D}^{*})$ of the Beltrami coefficients, its little subspace is defined as
[…]
The Teichmüller space of the little Zygmund smooth class is defined by $T^{Z}_{0}=\pi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$. We remark that the usage of naught 0 in this section is different from [10, 15].
We see that ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ is a closed subspace of ${\rm Bel}^{Z}(\mathbb{D}^{*})$. Indeed, assuming that a sequence $\{\mu_{k}\}_{k\in\mathbb{N}}$ in ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ and $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ are given so that $\|\mu_{k}-\mu\|_{Z}\to 0$ as $k\to\infty$, we can show that $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. For every $\varepsilon>0$, we choose some $k_{0}\in\mathbb{N}$ such that $\|\mu_{k_{0}}-\mu\|_{Z}<\varepsilon$. Since $\mu_{k_{0}}\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, there exists some $t_{0}>0$ such that
[…]
@@
uniformly as $t\to 0$. We can show that $(f^{\mu})^{\prime}$ satisfies the little Zygmund condition if and only if $\mu$ belongs to ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. Indeed, this is verified by examining the proof of [14, Theorem 1.1] with the fact that $(f_{\mu})^{\prime}|_{\mathbb{S}}$ satisfies the little Zygmund condition defined in a similar way to the case of self-diffeomorphisms of $\mathbb{S}$ if and only if
@@
To see that $\Phi:{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})\to A^{Z}_{0}(\mathbb{D})$ is a holomorphic split submersion, we verify that for every $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, the local holomorphic map $s_{\mu}:V_{\varphi}\to{\rm Bel}^{Z}(\mathbb{D}^{*})$ given in Theorem 2.4 maps $A^{Z}_{0}(\mathbb{D})\cap V_{\varphi}$ into ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. We first choose $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric on $\mathbb{D}^{*}$. By examining the proof of Lemma 2.3, we can choose such $\nu$ in ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$. Then, as in the proof of [10, Lemma 7.5], we have
@@
for every $\varphi^{\prime}\in V_{\varphi}$ and every $z^{*}=1/\bar{z}\in\mathbb{D}^{*}$. If $\varphi^{\prime}\in A^{Z}_{0}(\mathbb{D})$ in addition, then $\varphi^{\prime}-\varphi\in A^{Z}_{0}(\mathbb{D})$, and $\nu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ imply $s_{\nu}(\varphi^{\prime})\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$.
For the arbitrary $\mu\in{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$, we apply the right translations $r_{\mu}$ and $r_{\nu}$, which are biholomorphic automorphisms of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ by Lemma 2.2, and in fact of ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ as we can see by the proof of Proposition 2.1. Then, $s_{\mu}=r_{\mu}^{-1}\circ r_{\nu}\circ s_{\nu}$ is a local holomorphic right inverse of $\Phi$ defined on $V_{\varphi}\cap A^{Z}_{0}(\mathbb{D})$ satisfying $s_{\mu}\circ\Phi(\mu)=\mu$. ∎
Corollary 4.2.
@@
The claim $\Phi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))=A^{Z}_{0}(\mathbb{D})\cap\Phi({\rm Bel}(\mathbb{D}^{*}))$ is seen by the proof of Theorem 4.1 because $s_{\mu}(A^{Z}_{0}(\mathbb{D})\cap V_{\varphi})\subset{\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ implies $A^{Z}_{0}(\mathbb{D})\cap V_{\varphi}\subset\Phi({\rm Bel}^{Z}_{0}(\mathbb{D}^{*}))$. The other claims directly follows from Theorem 4.1. ∎
@@
The former claim is proved similarly to the case of the Schwarzian derivative map $\Phi$ as in the proof of Theorem 4.1. For the latter claim, we use the holomorphic map $\Lambda:B^{Z}(\mathbb{D})\to A^{Z}(\mathbb{D})$ with $\Lambda\circ\Psi=\Phi$, which is handled in Lemma 3.2. Since $\Phi$ is a holomorphic split submersion on ${\rm Bel}^{Z}_{0}(\mathbb{D}^{*})$ by Theorem 4.1, we see that so is $\Psi$. ∎